# Multidimensionality of the Hubble tension: the roles of $\Omega_{m}$ and $\omega_{c}$

### v1

@@ VI Conclusions
We have argued that, if both $\Omega_{m}$ and $\omega_{b}$ are calibrated (respectively by BAO and/or uncalibrated SNeIa, and by BBN considerations), then an increase in $H_{0}$ must necessarily be accompanied by an increase in $\omega_{c}$. Under the assumption that the primordial power spectrum of scalar fluctuations remains $\Lambda$CDM-like, an increase in $\omega_{c}$ implies, via the increase in $\omega_{m}$ (since $\omega_{b}$ is calibrated), implies an increase in the clustering parameter $S_{8}$, worsening the mild $S_{8}$ discrepancy. These shifts are not a consequence of the effects of new physics at early times, but follow solely from late-time expansion history constraints. Therefore, successful early-time new physics models must be able to accommodate these changes (which, in the case of $\omega_{c}$, are actually welcome as they help reduce the excess early ISW power typically associated to these models). As emphasized in the recent Ref. Poulin et al. (2024) (see also Refs. Clark et al. (2023); Vagnozzi (2023)), additional late-time new physics may help in this sense by weakening the constraints on $\Omega_{m}$ and/or $\sigma_{8}$. Although realizing a successful combination of early-plus-late new physics remains non-trivial, our work and Ref. Poulin et al. (2024) (as well as the earlier attempt in Ref. Toda et al. (2024)) emphasize the key role played by $\Omega_{m}$ in building such a combination. It is worth noting that, if the recent DESI BAO data are taken at face value, this task becomes somewhat easier.
@@
Our work has also underscored the crucial importance of a consistent $\Omega_{m}$ calibration. At present, different late-time probes of $\Omega_{m}$ are in mild disagreement between each other, with recent SNeIa catalogs pushing towards increasingly high values of $\Omega_{m}$: focusing only on the Pantheon versus PantheonPlus, we have explicitly demonstrated the effects of a different $\Omega_{m}$ calibration in exacerbating the aforementioned shifts in $\omega_{c}$ and $S_{8}$. Moving forward, we therefore believe that obtaining a reliable calibration of $\Omega_{m}$ which is as model-independent as possible should be a key priority in the cosmology community, in order to nail down one of the key actors in the Hubble tension play. We have argued that gas mass fraction measurements from galaxy clusters could be a promising probe in this sense, and we encourage the broader cosmology community to explore these and other less widely used probes. As we have officially entered the era of Stage IV cosmology a solution to the Hubble tension still eludes us, we note that the shifts in cosmological parameters which we have discussed (and which we have argued unavoidably accompany successful resolutions) inevitably leave their signatures in other cosmological observables, which current and upcoming cosmological surveys will hopefully be poised to detect Amendola et al. (2018); Abazajian et al. (2016); Aghamousa et al. (2016); Ade et al. (2019); Abitbol et al. (2019).

### v2

@@ VI Conclusions
We have argued that, if both $\Omega_{m}$ and $\omega_{b}$ are calibrated (respectively by BAO and/or uncalibrated SNeIa, and by BBN considerations), then an increase in $H_{0}$ must necessarily be accompanied by an increase in $\omega_{c}$. Under the assumption that the primordial power spectrum of scalar fluctuations remains $\Lambda$CDM-like, an increase in $\omega_{c}$, via the associated increase in $\omega_{m}$ (since $\omega_{b}$ is calibrated), implies an increase in the clustering parameter $S_{8}$, worsening the mild $S_{8}$ discrepancy. These shifts are not a consequence of the effects of new physics at early times, but follow solely from late-time expansion history constraints. Therefore, successful early-time new physics models must be able to accommodate these changes (which, in the case of $\omega_{c}$, are actually welcome as they help reduce the excess early ISW power typically associated to these models). As emphasized in the recent Ref. Poulin et al. (2024) (see also Refs. Clark et al. (2023); Vagnozzi (2023)), additional late-time new physics may help in this sense by weakening the constraints on $\Omega_{m}$ and/or $\sigma_{8}$. Although realizing a successful combination of early-plus-late new physics remains non-trivial, our work and Ref. Poulin et al. (2024) (as well as the earlier attempt in Ref. Toda et al. (2024)) emphasize the key role played by $\Omega_{m}$ in building such a combination. It is worth noting that, if the recent DESI BAO data are taken at face value, this task becomes somewhat easier.
@@
Our work has also underscored the crucial importance of a consistent $\Omega_{m}$ calibration. At present, different late-time probes of $\Omega_{m}$ are in mild disagreement between each other, with recent SNeIa catalogs pushing towards increasingly high values of $\Omega_{m}$: focusing only on the Pantheon versus PantheonPlus samples, we have explicitly demonstrated the effects of a different $\Omega_{m}$ calibration in exacerbating the aforementioned shifts in $\omega_{c}$ and $S_{8}$. Moving forward, we therefore believe that obtaining a reliable calibration of $\Omega_{m}$ which is as model-independent as possible should be a key priority in the cosmology community (see also Ref. Baryakhtar et al. (2024)), in order to nail down one of the key actors in the Hubble tension play. We have argued that gas mass fraction measurements from galaxy clusters could be a promising probe in this sense, and we encourage the broader cosmology community to explore these and other less widely used probes. As we have officially entered the era of Stage IV cosmology and a solution to the Hubble tension continues to elude us, we note that the shifts in cosmological parameters which we have discussed (and which we have argued unavoidably accompany successful resolutions) inevitably leave their signatures in other cosmological observables, which current and upcoming cosmological surveys will hopefully be poised to detect Amendola et al. (2018); Abazajian et al. (2016); Aghamousa et al. (2016); Ade et al. (2019); Abitbol et al. (2019).